OpenStudy (anonymous):

Find the induvidual equations of the lines represented by 3x^2+8xy-3y^2+2x-4y-1=0

OpenStudy (tkhunny):

An easy way might just be to find two points on each line. Try x = 0 and see what y-values are produced. Try x = 1 and see what y-values are produced. You might be almost done, after that.

OpenStudy (anonymous):

@tkhunny how does that help ?

OpenStudy (tkhunny):

How do two points help you determine a line? x = 0 leads to y = 1/4 -- There is a point that must be on both lines. You can try y = 0 leading to x = 1/2, but now, with only two points, we are going to get only one line. Forget we did y = 0. Try x = 1 and see what happens.

OpenStudy (anonymous):

wouldnt you get \[-3y^2-4y-1=0\]

OpenStudy (anonymous):

@tkhunny with the roots -1 and -1/3??

OpenStudy (anonymous):

@tkhunny aha I see your approach. good

OpenStudy (anonymous):

so, the y-intercepts of the two lines are: -1 and -1/3

OpenStudy (anonymous):

@Yahoo! do you follow?

OpenStudy (anonymous):

Yeah..

OpenStudy (anonymous):

there you go, now all you need is the slopes

OpenStudy (anonymous):

hint.. since the co-efficients of x^2 and y^2 are same, I think the two lines are perpendicular!

OpenStudy (anonymous):

Nice tkhunny :)

OpenStudy (tkhunny):

If it's easy, it's easy. No sense beating it to death. :-) Of course, we DID have to know the conic was degenerate, first.

OpenStudy (shubhamsrg):

@tkhunny can;'t it represent 2 points from 2 different lines ?

OpenStudy (shubhamsrg):

I;d treat it as a quadratic eqn in y and solve for y

OpenStudy (tkhunny):

@shubhamsrg That is a good point, and maybe I didn't stress that enough. I tried to pick the easiest possible values. In this case, it was clear enough that either x = 0 or y = 0 would result in a point on both lines.